# AnyCharV: Bootstrap Controllable Character Video Generation with Fine-to-Coarse Guidance

### v1

@@ 4.1 Experimental Setting
We build a character video dataset called CharVG to train our proposed model, which contains 5,482 videos from the Internet with various characters. These videos range from 7 to 47 seconds in length. The ReferenceNet and denoising UNet are both initialized from Stable Diffusion 1.5 (Rombach et al., 2022), where the motion module in the denoising UNet is initialized from AnimateDiff (Guo et al., 2024). The pose guider is initialized from ControlNet (Zhang et al., 2023a). We conduct both qualitative and quantitative evaluation for our method. For quantitative evaluation, we collect 10 character images and 10 target driving videos from the internet, then generate videos with every image-video pair, which results in 100 evaluation videos. We adopt FVD (Unterthiner et al., 2018), Dover++ (Wu et al., 2023a), and CLIP Image Score (Radford et al., 2021) to evaluate the generation quality. Please refer to Section A for detailed experimental settings.
@@
| w/o mask aug. | 2719.46 | 51.09 | 72.68 |
| w/o self-boosting | 2784.62 | 51.24 | 72.41 |
| AnyCharV (ours) | 2582.59 | 52.20 | 73.15 |
@@ 4.2 Comparison
We compare our approach with three recent state-of-the-art methods, including Make-A-Protagonist (Zhao et al., 2023), Viggle (Viggle, 2024), and MIMO (Men et al., 2024).
@@
As shown in Figure 3, Make-a-Protagonist performs worst as for lacking effective spatial and temporal modeling. Notably, our AnyCharV can preserve more detailed appearance and avoid lots of artifacts than Viggle and MIMO, especially looking at the generated arms and hands in Figure 3. Moreover, our approach can handle the complex human-object interactions very well, e.g., playing basketball, which can not be done with Make-a-Protagonist and Viggle. These results strongly affirm the effectiveness and robustness of our proposed AnyCharV.
@@
We show the results in Table 1. Our AnyCharV surpasses open-sourced Make-A-Protagonist by a significant margin, with 1.7% FVD, 51.2% DOVER++, and 2.49% CLIP Image Score improvements. Notably, AnyCharV performs better than MIMO by 8.08% regarding FVD, indicting the high reality and diversity of our method. Meanwhile, our proposed AnyCharV outperforms Viggle by 2.43% in terms of DOVER++ score, demonstrating the high aesthetic and technical quality of our generated videos. Considering Viggle and MIMO are both closed-source industrial products, our AnyCharV shows great effectiveness and robustness comparing with them. For inference cost, Make-A-Protagonist, Viggle, MIMO, and AnyCharV take 140, 2, 8, and 5 minutes to generate a 5 seconds 24 FPS video with resolution , respectively, further denoting the high efficiency of our approach.
@@
We conduct human evaluations to further evaluate our approach, comparing with three state-of-the-art methods. We gather 750 answers from 25 independent human raters, evaluating the identity preservation of the reference character, the motion consistency between the generated video and the target driving video, and the scene similarity and interaction. The average ranking is computed and shown in Table 2. It shows that our AnyCharV significantly outperforms the open-source Make-A-Protagonist and performs on par with the closed-source models Viggle and MIMO, demonstrating the effectiveness of our model.
@@ 4.3 Ablation Studies
We further do some ablation studies for our proposed framework, with deep analysis on self-boosting strategy, mask augmentation, and the choices of mask types.
@@
Mask Augmentation.
AnyCharV augments the segmentation mask during the first stage to reduce the negative effect caused by the fine mask shape. The qualitative and quantitative results shown in Table 3 and Figure 4 demonstrate the necessity of our design. Without mask augmentation, the appearance of the reference character is disturbed greatly.
@@
We conduct ablation on mask types where we use different masks during stage 1 and stage 2. In the first training stage, the model composes the reference character and the target video scene within the accurate region, where the fine segmentation mask is preferred, indicated by the great spatial information loss for both character and background scene, produced by lines ‘Box & Seg.’ and ‘Box & Box’ in Figure 5 and Table 4. In the second training stage, the model is expected to better preserve the details of the reference character. In this case, we use more loose mask constrain, i.e., coarse bounding box mask, to guide the character generation, eliminating the adverse effect caused by the fine segmentation mask shape. Such performance improvement can be validated in Table 4 and Figure 5.
@@ 4.4 Application
Our AnyCharV can also be used to generate a video with a reference image generated by text-to-image (T2I) models, e.g., FLUX (Black Forest, 2023), driving by a target video generated by text-to-video (T2V) models, e.g., HunyuanVideo (Kong et al., 2024). As shown in Figure 6, we first utilize FLUX to generate a reference image and HunyuanVideo to synthesize a target video, then the generated target video is used to drive the synthesized reference character. These results clearly the strength and flexibility of our AnyCharV, proving its versatility.
@@ 5 Conclusion
In this work, we introduce a novel framework AnyCharV with fine-to-coarse guidance for controllable character video generation under a two-stage training strategy. The self-supervised composition strategy with fine mask guidance in the first stage basically learn to drive a reference image with a target driving video to guarantee the motion correctness and target scene maintenance. Self-boosting training is then carried out via building interactions between the reference and target characters with coarse mask guidance, where the detailed identity of the reference character can be better preserved. AnyCharV clearly beats state-of-the-art open-source models and performs just as well as leading closed-source industrial products. Most importantly, AnyCharV can be used for images and videos created by T2I and T2V models, showing its strong ability to generalize.
@@ A.1 Dataset
We build a character video dataset called CharVG to train our proposed model, which contains 5,482 videos from the internet with various characters. These videos range from 7 to 47 seconds in length. We utilize DWPose (Yang et al., 2023) to extract the whole-body poses from these videos, and the estimated poses are rendered as pose skeleton image sequences. Meanwhile, we use YOLOv8 (Reis et al., 2023) to obtain the segmentation and bounding box masks for the characters presented in these videos.
@@ A.2 Implementation Details
The ReferenceNet and denoising UNet are both initialized from Stable Diffusion 1.5 (Rombach et al., 2022), where the motion module in the denoising UNet is initialized from AnimateDiff (Guo et al., 2024). The pose guider is initialized from ControlNet (Zhang et al., 2023a). In our training, firstly, we train the ReferenceNet, denoising UNet without motion module, and pose guider with individual frames from videos for 50,000 steps to learn the spatial information. Secondly, we perform self-boosted training on only denoising UNet without motion module with individual video frames for 3,000 steps to help the model eliminate the bad influence of the mask shape. Above two training processes are both conducted with resolution and batch size 64. Thirdly, we only train the motion module in the denoising UNet with 24-frame video clips with resolution for 10,000 steps to improve the temporal consistency. Last, the self-boosted training strategy is performed only on the denoising UNet with resolution for 10,000 steps to better preserve the identity of the reference character. Above two training processes are both conducted with batch size 8. All above are trained using learning rate 1e-5. During inference, the long video is generated sequentially with several short clips and temporally aggregated following (Tseng et al., 2023). We use a DDIM (Song et al., 2020) scheduler for 30 denoising steps with classifier-free guidance (Ho & Salimans, 2021) as 3.0. All of our experiments are finished on 8 NVIDIA H800 GPUs using PyTorch (Paszke et al., 2019). The overall training process takes about 3 days. Our model takes 5 minutes to generate a 5 seconds 24FPS video with resolution on a single NVIDIA H800 GPU.
@@ A.3 Evaluation Metrics
We adopt several metrics to evaluate the generation quality. 1) FVD (Unterthiner et al., 2018) score is a widely used metric for assessing the generated videos. We compute FVD score between our generated videos and 1,000 real character videos from our dataset. 2) Dover++ (Wu et al., 2023a) score is a video quality assessment metric from both aesthetic and technical perspective, demonstrating the overall quality of the generated videos. 3) CLIP Image Score (Radford et al., 2021) is used to evaluate the similarity between the generated video and the reference character, validating the model capability for preserving the identity.
@@ A.4 Comparison Methods
Make-A-protagonist (Zhao et al., 2023) memorizes the visual and motion information of the target video by fine-tuning the video generation model and guiding the generation of the desired output through masks and reference images during the inference stage.
Viggle (Viggle, 2024) performs rapid 3D modeling of the reference image and manipulates it according to the pose sequence of the input video, thereby accomplishing tasks such as character generation.
MIMO (Men et al., 2024) decomposes the character’s identity, pose, and scene to facilitate video synthesis.
@@ Appendix B Additional Experimental Results
B.1 Qualitative Results
B.2 User Interface
@@ Appendix C Limitations and Future Works
While our AnyCharV model demonstrates impressive results in controllable character video generation, it may struggle when tasked with inferring the back view of a character from a front-facing reference image. Looking ahead, we plan to enhance the generalization capabilities of AnyCharV by integrating a more robust video generator and introducing more robust controls for open-world scenarios (Wang et al., 2024d; Che et al., 2024; Wang et al., 2024e).

### v2

@@ 4.1 Experimental Setting
We build a character video dataset called CharVG to train our proposed model, which contains 9,055 videos from the Internet with various characters. These videos range from 7 to 59 seconds in length. The ReferenceNet and denoising UNet are both initialized from Stable Diffusion 1.5 [48], where the motion module in the denoising UNet is initialized from AnimateDiff [49]. The pose guider is initialized from ControlNet [50]. We conduct both qualitative and quantitative evaluation for our method. For quantitative evaluation, we collect 20 character images and 10 target driving videos from the internet, then generate videos with every image-video pair, which results in 200 evaluation videos. We adopt FVD [51], Dover++ [52], and CLIP Image Score [45] to evaluate the generation quality. Please refer to Section A for detailed experimental settings.
@@
MAP [16] | 671.44 | 34.91 | 66.03 |
Viggle [15] | 638.92 | 55.47 | 67.97 |
MIMO [14] | 647.12 | 54.89 | 68.01 |
VACE [18] | 658.56 | 55.26 | 67.93 |
AnyCharV (ours) | 622.01 | 57.07 | 69.14 |
Method | Identity | Motion | Scene |
MAP [16] | 4.780.05 | 4.910.07 | 4.990.05 |
Viggle [15] | 2.040.24 | 2.200.33 | 2.040.31 |
MIMO [14] | 2.390.26 | 2.120.26 | 2.300.27 |
VACE [18] | 3.900.34 | 3.750.31 | 3.680.28 |
AnyCharV (ours) | 1.890.24 | 2.020.37 | 1.990.27 |
Method | FVD | DOVER++ | CLIP-I |
w/o mask inp. | 680.31 | 54.09 | 66.12 |
w/o char. seg. | 678.51 | 55.30 | 67.97 |
w/o mask aug. | 646.03 | 55.58 | 68.72 |
w/o self-boost | 652.34 | 53.54 | 67.75 |
AnyCharV (ours) | 622.01 | 57.07 | 69.14 |
@@ 4.2 Comparison
We compare our approach with four recent state-of-the-art methods, including Make-A-Protagonist [16], Viggle [15], MIMO [14], and VACE [18].
@@
As shown in Figure 3, Make-a-Protagonist performs worst as for lacking effective spatial and temporal modeling. Notably, our AnyCharV can preserve more detailed appearance and avoid lots of artifacts than Viggle and MIMO, especially looking at the generated arms and hands in Figure 3. VACE can not preserve the identity of the reference character very well. Moreover, our approach can handle the complex human-object interactions very well, e.g., playing basketball, which can not be done with Make-a-Protagonist, Viggle, and VACE. These results strongly affirm the effectiveness and robustness of our proposed AnyCharV.
@@
We show the results in Table 2. Our AnyCharV surpasses the best open-sourced model VACE by a significant margin, with 5.55% FVD, 3.17% DOVER++, and 1.78% CLIP Image Score improvements. Moreover, AnyCharV performs much better than Viggle and MIMO, which are both closed-source industrial products. For example, AnyCharV outperforms MIMO by 3.88%, 3.97%, 1.66% in terms of FVD, DOVER++, and CLIP Image Score. These results clearly demonstrate the great effectiveness and robustness comparing with such closed-source industrial product. For inference cost, Make-A-Protagonist, Viggle, MIMO, VACE, and AnyCharV take 140, 2, 8, 10, and 5 minutes to generate a 5 seconds 24 FPS video with resolution , respectively, further denoting the high efficiency of our approach.
@@
We conduct human evaluations to further evaluate our approach, comparing with four SOTA methods. We gather 1,200 answers from 30 independent human raters, evaluating the identity preservation of the reference character, the motion consistency between the generated video and the target driving video, and the scene similarity and interaction. The average ranking is computed and shown in Table 2. This shows that our AnyCharV significantly outperforms the open-source Make-A-Protagonist and MACE as well as the closed-source models Viggle and MIMO, demonstrating the effectiveness of our model. More details of the metrics can be found in Section A.3.
@@ 4.3 Ablation Studies
Mask Input.
For both two stages, we provide the mask information to help the model locate the area of the character and further composition. From the results in from Table 4 and Figure 5, we find that without the mask input, the model fails to compose the source character and target scene. These results also demonstrate that masking the character in the target video can not incorporate fine segmentation information in the first stage to guide the composition.
Character Segmentation.
During training, we segment the character out for all reference images to avoid introducing redundant context information which can degrade the quality of the generated videos. From Table 4 and Figure 5, we can observe that the unnecessary background in the reference character will dramatically degrade the quality of the generated character. These results validate the necessity for the segmentation of characters.
Mask Augmentation.
AnyCharV augments the segmentation mask during the first stage to reduce the negative effect caused by the fine mask shape. The qualitative and quantitative results shown in Table 4 and Figure 5 demonstrate the effectiveness of our design. Without mask augmentation, the appearance of the reference character is disturbed greatly.
@@
We conduct ablation on different mask types during two stages. In the first training stage, the model composes the reference character and the target video scene within the accurate region, where the fine segmentation mask is preferred, indicated by the great spatial information loss for both character and background scene, produced by lines ‘Box & Seg.’ and ‘Box & Box’ in Figure 5 and Table 4. In the second training stage, the model is expected to better preserve the details of the reference character. In this case, we use more loose mask constrain, i.e., coarse bounding box mask, to guide the character generation, eliminating the adverse effect caused by the fine segmentation mask shape. Such performance improvement can be validated in Table 4 and Figure 5.
@@ 4.4 Application
Our AnyCharV can also be used to generate a video with a reference image generated by text-to-image (T2I) models, e.g., FLUX [53], driving by a target video generated by text-to-video (T2V) models, e.g., HunyuanVideo [54]. As shown in Figure 6, we first utilize FLUX to generate a reference image and HunyuanVideo to synthesize a target video, then the generated target video is used to drive the synthesized reference character. These results clearly indicate the strength and flexibility of our AnyCharV, proving its versatility.
@@ 5 Conclusion
In this work, we introduce a novel framework AnyCharV with fine-to-coarse guidance for controllable character video generation under a two-stage training strategy. The self-supervised composition strategy with fine mask guidance in the first stage basically learn to drive a reference image with a target driving video to guarantee the motion correctness and target scene maintenance. Self-boosting training is then carried out via building interactions between the reference and target characters with coarse mask guidance, where the detailed identity of the reference character can be better preserved. AnyCharV clearly beats both SOTA open-source models and leading closed-source industrial products. Most importantly, AnyCharV can be used for images and videos created by T2I and T2V models, showing its strong ability to generalize.
@@ A.1 Dataset
We build a character video dataset called CharVG to train our proposed model, which contains 9,055 videos from the internet with various characters. These videos range from 7 to 59 seconds in length. We utilize DWPose [46] to extract the whole-body poses from these videos, and the estimated poses are rendered as pose skeleton image sequences. Meanwhile, we use YOLOv8 [55] and SAM2 [44] to obtain the segmentation and bounding box masks for the characters presented in these videos.
@@ A.2 Implementation Details
The ReferenceNet and denoising UNet are both initialized from Stable Diffusion 1.5 [48], where the motion module in the denoising UNet is initialized from AnimateDiff [49]. The pose guider is initialized from ControlNet [50]. In our training, firstly, we train the ReferenceNet, denoising UNet without motion module, and pose guider with individual frames from videos for 50,000 steps to learn the spatial information. Secondly, we perform self-boosted training on only denoising UNet without motion module with individual video frames for 3,000 steps to help the model eliminate the bad influence of the mask shape. Above two training processes are both conducted with resolution and batch size 64. Thirdly, we only train the motion module in the denoising UNet with 24-frame video clips with resolution for 10,000 steps to improve the temporal consistency. Last, the self-boosted training strategy is performed only on the denoising UNet with resolution for 10,000 steps to better preserve the identity of the reference character. Above two training processes are both conducted with batch size 8. All above are trained using learning rate 1e-5. During inference, the long video is generated sequentially with several short clips and temporally aggregated following [56]. We use a DDIM [30] scheduler for 30 denoising steps with classifier-free guidance [57] as 3.0. All of our experiments are finished on 8 NVIDIA H800 GPUs using PyTorch [58]. The overall training process takes about 3 days. Our model takes 5 minutes to generate a 5 seconds 24FPS video with resolution on a single NVIDIA H800 GPU.
@@ A.3 Evaluation Metrics
Quantitative Evaluation.
We adopt three metrics to evaluate the generation quality:
FVD [51] score is a widely used metric for assessing the generated videos. We compute FVD score between our generated videos and 1,000 real character videos from our dataset.
Dover++ [52] score is a video quality assessment metric from both aesthetic and technical perspective, demonstrating the overall quality of the generated videos.
CLIP Image Score [45] is used to evaluate the similarity between the generated video and the reference character, validating the model capability for preserving the identity.
User Study.
In user study, the quality of the generated video are assessed from three aspects:
Character identity preservation (Identity in Table 2): this metric is used to evaluate whether the reference character remains unchanged in the generated videos.
Motion consistency (Motion in Table 2): this metric measures the fluidity of the character’s motion and whether the motions match between the target video and the generated video.
Scene similarity/interaction (Scene in Table 2): this metric assesses the preservation of the background scene shown in the target video. Also paying attention to the interaction, as the object is also a part of the background scene.
We have tried our best to carry out the user study under a diverse scenario from two aspects:
Various samples: we generate various videos with different types of target scenes (sports, dancing, film, …) and different reference characters (real human, anime character, AI generated character, …).
Diverse human raters: the generated videos are evaluated by human raters with diverse backgrounds, including relevant researchers, financial practitioners, teachers, and so on.
@@ A.4 Comparison Methods
Make-A-protagonist [16] memorizes the visual and motion information of the target video by fine-tuning the video generation model and guiding the generation of the desired output through masks and reference images during the inference stage.
Viggle [15] performs rapid 3D modeling of the reference image and manipulates it according to the pose sequence of the input video, thereby accomplishing tasks such as character generation.
MIMO [14] decomposes the character’s identity, pose, and scene to facilitate video synthesis. And it can be used for composing the character and scene.
VACE [18] learns various video generation and editing tasks in a unified manner, which can be applied for character replacement.
@@ Appendix B Additional Experimental Results
Synthesized Video Pairs.
We build 64,000 video pairs for training in the second stage. Here, we illustrate some synthesized video pairs in Figure 7. We could learn that the high-quality synthesized video pairs strongly support the effective learning in the second stage.
Qualitative Results.
User Interface.
@@
Demo Video.
We attach a demo video in the supplementary material, including three parts: 1) comparison with SOTA methods; 2) generated videos with our AnyCharV; 3) application: combining with Text-to-Image and Text-to-Video models. To be note that MIMO automatically intercepts the first half in the first comparison video and VACE automatically sample a subset of frames for generation.
@@ Appendix C Limitations and Future Works
While our AnyCharV model demonstrates impressive results in controllable character video generation, it may struggle when tasked with inferring the back view of a character from a front-facing reference image. Looking ahead, we plan to enhance the generalization capabilities of AnyCharV by integrating a more robust video generator and introducing more specialized controls for specific scenarios.
Appendix D Broader Impact
Flexibly controlling the generation of different characters is crucial in our practical life. We achieve composing an arbitrary character and an arbitrary target scene with our proposed two-stage framework AnyCharV. We strongly believe that our work will be beneficial for a number of AIGC applications, such as creative AI and film production. However, our model still suffers from generating the back view of the provided character. Please be careful to use our model when meeting such scenario.